let f(x)=2x+1 and g(x)=3x to the power of 2. find Fog
Answer by stanbon(48558)

let f(x)=2x+1 and g(x)=3x^2. find Fog
---------------
Fog = F[g(x)] = F[3x^2] = 2[3x^2]+1 = 6x^2+1
==============
Cheers,
Stan H.